OpenStudy (anonymous):

so what is the vertex of f(x)=3x^2-12x+21

OpenStudy (anonymous):

use: -b/2a, to find the x coordinate of the vertex, then once you find it, you can plug that value back into the orginal function and get the y coordiante of the vertex

OpenStudy (anonymous):

so the vertex would be (3,21)

OpenStudy (anonymous):

vertex would be(2,9)

OpenStudy (anonymous):

Good try i guess

OpenStudy (across):

Use calculus; things simplify tenfold:\[f'(x)=6x-12=0\implies x=2\]\[f(2)=3(2)^2-12(2)+21=9\]\[(2,9)\]and Lagrange is correct.